# RadioDiff-FS: Physics-Informed Manifold Alignment in Few-Shot Diffusion Models for High-Fidelity Radio Map Construction

### 19 Mar 2026 v1

@@ V-C Performance on SRM
Quantitative results on the SRM are reported in Table I. RadioDiff-FS achieves the best performance across all metrics, attaining an NMSE of $0.0049$ and a PSNR of $36.37$ dB. Relative to the vanilla RadioDiff baseline with NMSE $0.0121$, and PhyRMDM with NMSE $0.0100$, this corresponds to a clear reduction in reconstruction error, indicating that few-shot training alone is insufficient to reach high-fidelity multipath reconstruction without additional adaptation bias. RadioDiff-FS also yields the highest SSIM of $0.9752$, substantially exceeding RadioUNet, which suggests improved preservation of interaction-driven spatial structures such as diffraction edges and shadow boundaries. These trends are consistent with the qualitative results in Fig. 3. RadioUNet produces smooth attenuation fields that capture coarse pathloss gradients but blur sharp transitions near building silhouettes, which are critical in multipath-dominant settings. RME-GAN generates sharper textures, yet it exhibits intermittent artifacts in low-signal regions, where adversarial objectives may amplify visually plausible but physically inconsistent patterns. Diffusion-based baselines improve perceptual realism, but without an explicit cross-domain constraint they still under-represent boundary sharpness and local contrast in deep-shadow areas. In contrast, RadioDiff-FS better matches the ground truth both globally and locally: it preserves the large-scale attenuation skeleton while recovering fine-grained transition bands around occluded corridors, aligning with the intended main-path to multipath generalization behavior.
@@ V-D Performance on DRM
The DRM further introduce time-varying occlusions from vehicles, which produce localized attenuation spots and small-scale shadowing patterns along roads. As shown in Table II, most baselines degrade noticeably when moving from SRM to DRM. In particular, PhyRMDM suffers a sharp deterioration with NMSE increasing to $0.1113$, indicating that physics-alignment alone does not guarantee robustness when the target domain contains sparse, localized, and configuration-dependent perturbations under limited supervision. The vanilla RadioDiff baseline also degrades, reflecting that the pretrained generative prior may not correctly allocate capacity to vehicle-induced details when trained with few DRM samples. RadioDiff-FS remains substantially more stable in this regime, achieving an NMSE of $0.0121$, which is markedly lower than PhyRMDM and also improves upon RME-GAN. The qualitative comparison in Fig. 4 explains this stability. Vehicle occlusions manifest as compact attenuation patches whose locations are coupled to road geometry and line-of-sight corridors. Regression and diffusion baselines tend to smooth these patches into the background, leading to systematic underestimation of localized blockage. RadioDiff-FS more consistently recovers these attenuation signatures and their spatial extent, producing DRM reconstructions that are closer to the ground truth in both contrast and location. This behavior is particularly important for downstream tasks that depend on local reliability margins, where missing a small occlusion-induced dip can lead to incorrect link-quality ranking.
@@ V-E Ablation Study
We ablate the proposed direction-consistency loss (DCL) to quantify its contribution under both full fine-tuning and low-rank adaptation (LoRA). DCL is designed to constrain the feature displacement of generated maps to follow the dominant main-path to multipath shift while suppressing orthogonal distortions, which empirically translates to sharper boundaries and fewer spurious artifacts in low-signal regions. Table IV highlights the more challenging DRM setting. Removing DCL degrades performance consistently: under full fine-tuning, adding DCL reduces NMSE from $0.0163$ to $0.0121$ and improves SSIM from $0.9105$ to $0.9510$. The SSIM gain indicates that DCL primarily improves structural consistency rather than merely rescaling pixel intensities, matching the goal of preserving interaction-induced patterns. The improvement remains visible under LoRA, suggesting that DCL is complementary to parameter-efficient adaptation and helps concentrate limited update capacity on multipath-relevant corrections.
Similar trends are observed on SRM in Table III. Although SRM is less perturbed than DRM, DCL still yields consistent gains for both adaptation strategies. Taken together, these ablations confirm that the advantage of RadioDiff-FS does not stem solely from additional training steps, but from the proposed directional constraint that provides a principled and effective inductive bias for few-shot generalization from main-path structure to multipath detail.
@@
TABLE IV: Ablation Study on IRT4 With Car. Comparison of Full Fine-Tuning (Full FT) and LoRA strategies with and without Direction-Consistency Loss (DCL).

### 26 Mar 2026 v3

@@ V-C Performance on SRM
Quantitative results on the SRM are reported in Table I. RadioDiff-FS achieves the best performance across all metrics, attaining an NMSE of $0.0049$ and a PSNR of $36.37$ dB. Compared with the vanilla RadioDiff baseline at NMSE $0.0121$ and PhyRMDM at NMSE $0.0100$, RadioDiff-FS reduces the reconstruction error by $59.50\%$. This result indicates that few-shot training alone is insufficient for high-fidelity multipath reconstruction without additional adaptation guidance. RadioDiff-FS also yields the highest SSIM of $0.9752$, which substantially exceeds all other methods and suggests improved preservation of spatial structures such as diffraction edges and shadow boundaries. The one-shot variant of RadioDiff-FS also demonstrates competitive performance. With only a single MU-RM sample per scene for fine-tuning, RadioDiff-FS achieves the second-best SSIM of $0.9285$ and the second-best PSNR of $31.77$ dB, outperforming all baselines except the full RadioDiff-FS. This result validates that the directional consistency constraint provides a strong inductive bias even under extreme data scarcity, as it leverages the pretrained MP-RM prior to guide adaptation with minimal supervision.
These quantitative trends are consistent with the qualitative results in Fig. 3. RadioUNet produces smooth attenuation fields that capture coarse pathloss gradients but blur sharp transitions near building silhouettes. These transitions are critical in multipath-dominant settings. RME-GAN generates sharper textures, yet it exhibits intermittent artifacts in low-signal regions where adversarial objectives may amplify visually plausible but physically inconsistent patterns. Diffusion-based baselines improve perceptual realism. However, without an explicit cross-domain constraint, they still under-represent boundary sharpness and local contrast in deep-shadow areas. In contrast, RadioDiff-FS better matches the ground truth both globally and locally. It preserves the large-scale attenuation skeleton while recovering fine-grained transition bands around occluded corridors, aligning with the intended main-path to multipath generalization behavior.
@@ V-D Performance on DRM
The DRM further introduces time-varying occlusions from vehicles, which produce localized attenuation spots and small-scale shadowing patterns along roads. As shown in Table II, most baselines degrade noticeably when moving from SRM to DRM. PhyRMDM suffers a sharp deterioration with NMSE increasing to $0.1113$, indicating that physics-alignment alone does not guarantee robustness when the target domain contains sparse and configuration-dependent perturbations under limited supervision. The vanilla RadioDiff baseline also degrades significantly, reflecting that the pretrained generative prior may not correctly allocate capacity to vehicle-induced details when trained with few DRM samples. RadioDiff-FS remains substantially more stable in this regime, achieving an NMSE of $0.0121$ and an SSIM of $0.9510$, which are markedly better than all baselines. The one-shot variant again delivers strong results, attaining the second-best RMSE of $0.0296$ and the second-best SSIM of $0.8969$. This confirms that the directional consistency constraint is effective even with a single fine-tuning sample per scene.
The qualitative comparison in Fig. 4 explains this stability. Vehicle occlusions manifest as compact attenuation patches whose locations are coupled to road geometry and line-of-sight corridors. Regression and diffusion baselines tend to smooth these patches into the background, leading to systematic underestimation of localized blockage. RadioDiff-FS more consistently recovers these attenuation signatures and their spatial extent, producing DRM reconstructions that are closer to the ground truth in both contrast and location. This behavior is particularly important for downstream tasks that depend on local reliability margins, where missing a small occlusion-induced dip can lead to incorrect link-quality assessment.
@@ V-E Ablation Study
We ablate the proposed direction-consistency loss (DCL) to quantify its contribution under both full fine-tuning and low-rank adaptation (LoRA). The DCL constrains the feature displacement of generated maps to follow the dominant main-path to multipath shift while suppressing orthogonal distortions. This empirically translates to sharper boundaries and fewer spurious artifacts in low-signal regions. Table IV presents results on the more challenging DRM setting. Removing DCL degrades performance consistently. Under full fine-tuning, adding DCL reduces NMSE from $0.0163$ to $0.0121$ and improves SSIM from $0.9105$ to $0.9510$. The SSIM gain indicates that DCL primarily improves structural consistency rather than merely rescaling pixel intensities, which matches the goal of preserving interaction-induced patterns. The improvement remains visible under LoRA, where DCL reduces NMSE from $0.0259$ to $0.0195$ and improves SSIM from $0.8750$ to $0.8982$. This suggests that DCL is complementary to parameter-efficient adaptation and helps concentrate limited update capacity on multipath-relevant corrections.
Similar trends are observed on SRM in Table III. Although SRM is less challenging than DRM, DCL still yields consistent gains for both adaptation strategies. Under full fine-tuning, DCL improves PSNR from $35.79$ dB to $36.37$ dB and SSIM from $0.9715$ to $0.9752$. Under LoRA, the gains follow the same direction with PSNR improving from $31.88$ dB to $32.65$ dB. Taken together, these ablations confirm that the advantage of RadioDiff-FS does not stem solely from additional training steps. The proposed directional constraint provides a principled and effective inductive bias for few-shot generalization from main-path structure to multipath detail.
@@
TABLE IV: Ablation Study on IRT4 With Car. Comparison of full fine-tuning and LoRA strategies with and without direction-consistency loss (DCL).